# thanks: Equal contributorsthanks: Equal contributors

### v2

@@ A Taylor expansion of the spring energy around this state gives that
The numerical framework for the constrained optimization problem employs the fmincon toolbox to perform a sequential quadratic programming (SQP) algorithm in Matlab (R2023a). The eight optimization variables include the components of the vectors $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ and four rotation angles in array $\bm{\phi}$. These variables undergo an iterative update process, transforming the kirigami geometry from its initial guess to an optimal design that accomplishes the target properties and meets the nonlinear constraints. Given the non-convex nature of the objective functions and constrained sets, the optimized geometry obtained by each simulation is dependent upon the choice of the starting point. As illustrated by way of an example in Fig. 4A (main text), different initial guesses for the design can yield different optimized designs that achieve the specified overall deformation and target energy barrier. Also, in some of the cases of extreme shape change in Fig. 4A-B (main text), we are not able to achieve the target energy barrier, so we test several initial guesses to ensure that the deviation from the target is robust. We have also explored other objective functions consistent with minimizing the same objective, e.g., by replacing $|\cdot|^{2}$ terms with ||$\cdot$|| ⋅ | terms, and found that such modifications can sometimes lead to improvements in the algorithms ability to find designs that get closer to the overall objective. To avoid ad hoc procedures, we only present results for cases where the objective function and its fitting parameters are consistent across all optimized designs being compared.
@@
S3.B Supporting information on examples
@@
Planar kirigami found in the literature are typically monolithic, and feature panels connected via elastic hinges. A common way to fabricate such metamaterials is by laser cutting rubber. While this strategy works well for metamaterials featuring low energy modes of deformation, it does not work well for bistable metamaterials; in fact, there is a limit to how thick of a rubber sheet one can laser cut, and thin sheets can lead to low out-of-plane stiffness and, potentially, to loss of bistability. Thus, we adopt a strategy previously employed by Wu and Pasini [53], where thick monolithic metamaterials are 3D printed out of thermoplastic polyurethane (TPU). In particular, we use a TPU-95A filament on an Ultimaker 2+ printer. The steps we follow to design and fabricate these specimens are reported below. We let $\ell_{1}^{R}=|\bm{\ell}_{1}^{R}|$ and $\ell_{2}^{R}=|\bm{\ell}_{2}^{R}|$ characterize the lengths of the physical specimens.
@@
To avoid dealing with hinge energy, which can prevent patterns from exhibiting bistability, we resort to a design for kirigami where panels are connected via actual pin joints. To try and create specimens that deform in plane, we consider a skeletal version of each panel, i.e., we replace each panel by an assembly of eight bars and five hinges. To make sure that deformations are not localized at the hinges, we fabricate hinge regions out of stiff Nylon and bars out of softer TPU-95A. Each panel is fabricated in a single shot, leveraging the dual-extrusion nature of our Ultimaker 3 printer. The steps we follow to design and fabricate these bi-material specimens are reported below.
@@
Our final physical incarnation of the theoretically-designed bistable patterns features thin mono-material panels laser cut out of PETG and assembled with push-in rivets. These specimens are designed to bend out of plane as they transition between stable states, and the steps we follow to design and fabricate them are reported below.

### v3

@@ A Taylor expansion of the spring energy around this state gives that
Figure S3: Flowchart of numerical aspects to the optimization framework.
@@
The numerical framework for the constrained optimization problem employs the fmincon toolbox to perform a sequential quadratic programming (SQP) algorithm in Matlab (R2023a). A flowchart of the procedure is illustrated in Fig S3. To enable the optimization, we prescribe the Bravais lattice vectors $\bm{\ell}$ for two stable states and constrain the design vectors $\mathbf{s}_{3},\ldots,\mathbf{v}$ according to Eq. (S22). The eight remaining variables — the design vectors $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ and four rotation angles in the array $\bm{\phi}$ — are then optimized based on an assigned objective function $f_{\text{obj}}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})$ (for instance Eq. (S55)) and the inequality constraints $\mathbf{g}_{\text{ineq}}(\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi})$ in Eq. (S24). Beginning with an initial guess $\mathbf{x}_{0}=(\mathbf{s}_{1}^{0},\mathbf{s}_{2}^{0},\bm{\phi}^{0})$, these variables undergo iterative update $\mathbf{x}_{i}=(\mathbf{s}_{1}^{i},\mathbf{s}_{2}^{i},\bm{\phi}^{i})$ using the SQP algorithm until a local minimum of the energy is achieved. This procedure transforms the kirigami geometry from its initial guess $\mathbf{x}_{0}$ to an optimal design $\mathbf{x}_{\text{opt}}=(\mathbf{s}_{1}^{\text{opt}},\mathbf{s}_{2}^{\text{opt}},\bm{\phi}^{\text{opt}})$ that accomplishes the target properties and meets the nonlinear constraints.
Given the non-convex nature of the objective functions and constrained sets, the optimized geometry obtained by each simulation is dependent upon the choice of the starting point $\mathbf{x}_{0}$. As illustrated by way of an example in Fig. 4A (main text), different initial guesses for the design can yield different optimized designs that achieve the specified overall deformation and target energy barrier. Also, in some of the cases of extreme shape change in Fig. 4A-B (main text), we are not able to achieve the target energy barrier, so we test several initial guesses to ensure that the deviation from the target is robust. We have also explored other objective functions consistent with minimizing the same objective, e.g., by replacing $|\cdot|^{2}$ terms with ||$\cdot$|| ⋅ | terms, and found that such modifications can sometimes lead to improvements in the algorithms ability to find designs that get closer to the overall objective. To avoid ad hoc procedures, we only present results for cases where the objective function and its fitting parameters are consistent across all optimized designs being compared.
@@
Figure S6: Two marching algorithm procedures and the corresponding heterogeneous configuration results (Letters ’S’ in red and ’E’ in blue denote the starting and ending cell). (A) From the lower-left corner to the upper-right (algorithm shown in the previous section). (B) Spiral marching from the center towards the right-bottom corner.
S3.B Robustness to marching strategy
This section discusses the robustness of examples of heterogeneous shape change to the choice of marching strategy. One strategy entails starting from lower-left unit cell and marching in a zig-zag fashion as depicted in Fig. S6-A. Another involves initializing the algorithm from the central cell and marching by spiraling out as shown in Fig. S6-B. In the case of the square-to-bowtie transformation, we have investigated both ways of marching, keeping all other details of the procedure fixed. The right part of Fig. S6 shows the examples that emerge from these two ways marching. The design variables do change slightly from cell-to-cell, as indicated by the table below. However, the overall designs are qualitatively quite similar to each other. This result suggests that the marching algorithm is robust to small changes in the way the algorithm is initialized. In other words, the design space for heterogeneous shape change seems to be highly constrained rather than highly degenerate.
Design variables $\underset{\{i,j\}}{\max}\frac{|\mathbf{x}^{A}(i,j)-\mathbf{x}^{B}(i,j)|}{|\mathbf{x}^{A}(i,j)|}$
$\mathbf{x}=[\mathbf{s}_{1},\mathbf{s}_{2},\bm{\phi}]$ $1.057\%$
S3.C Supporting information on examples
  • Refer to caption
@@
Planar kirigami found in the literature are typically monolithic, and feature panels connected via elastic hinges. A common way to fabricate such metamaterials is by laser cutting rubber. While this strategy works well for metamaterials featuring low energy modes of deformation, it does not work well for bistable metamaterials; in fact, there is a limit to how thick of a rubber sheet one can laser cut, and thin sheets can lead to low out-of-plane stiffness and, potentially, to loss of bistability. Thus, we adopt a strategy previously employed by Wu and Pasini [58], where thick monolithic metamaterials are 3D printed out of thermoplastic polyurethane (TPU). In particular, we use a TPU-95A filament on an Ultimaker 2+ printer, since this is the softest material of the Ultimaker material palette. The steps we follow to design and fabricate these specimens are reported below. We let $\ell_{1}^{R}=|\bm{\ell}_{1}^{R}|$ and $\ell_{2}^{R}=|\bm{\ell}_{2}^{R}|$ characterize the lengths of the physical specimens.
@@
To avoid dealing with hinge energy, which can prevent patterns from exhibiting bistability, we resort to a design for kirigami where panels are connected via actual pin joints. To try and create specimens that deform in plane, we consider relatively-thick specimens where each panel is skeletal in nature, i.e., we replace each panel by an assembly of eight bars and five hinges. To make sure that deformations are not localized at the hinges and to avoid excessive dissipation caused by pins rotating against soft materials, we fabricate hinge regions out of stiff Nylon and bars out of softer TPU-95A. Each panel is fabricated in a single shot, leveraging the dual-extrusion nature of our Ultimaker 3 printer. The steps we follow to design and fabricate these bi-material specimens are reported below.
@@
Our final physical incarnation of the theoretically-designed bistable patterns features thin mono-material panels laser cut out of PETG and assembled with push-in rivets. We choose PETG since it provides the right amount of compliance with respect to more classical laser cuttable polymers like acrylic; the thickness is chosen to be small to avoid high stresses that could lead to breaking of the panels near the relatively-weak hinge regions. These specimens are designed to bend out of plane as they transition between stable states, and the steps we follow to design and fabricate them are reported below.